OpenStudy (anonymous):

what is the equation of direct variation, given that y varies directly with x. y is 15 when x is 30

OpenStudy (tkhunny):

"y varies directly with x" MEANS y = kx. Just write it down when you hear it..

jimthompson5910 (jim_thompson5910):

y = kx 15 = k*30 ... plug in x = 30 and y = 15 solve for k and tell me what you get

jimthompson5910 (jim_thompson5910):

are you able to find k at all?

OpenStudy (anonymous):

15+30=-15 ?

jimthompson5910 (jim_thompson5910):

k*30 means k times 30

jimthompson5910 (jim_thompson5910):

how do you undo multiplication?

OpenStudy (anonymous):

30-15=15

jimthompson5910 (jim_thompson5910):

division is the opposite of multiplication, agreed?

OpenStudy (anonymous):

oohhhh 30 /15=2

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

but you flipped them around

OpenStudy (anonymous):

0.5

jimthompson5910 (jim_thompson5910):

or 1/2

jimthompson5910 (jim_thompson5910):

so k = 1/2

jimthompson5910 (jim_thompson5910):

and here's the full step by step picture y = kx 15 = k*30 15/30 = k*30/30 15/30 = k*1 1/2 = k k = 1/2 ----------------------------------------- \[\Large y = kx\] \[\Large y = \frac{1}{2}x\] ----------------------------------------- So the answer is \[\Large y = \frac{1}{2}x\]

OpenStudy (anonymous):

okay thanks i get it alot better now!(:

jimthompson5910 (jim_thompson5910):

that's great, glad it's all clicking now, yw

OpenStudy (anonymous):

@jim_thompson5910 strikes again lol By the way he is one of the best teachers :)

OpenStudy (anonymous):

lol im a slow thinker that all!

jimthompson5910 (jim_thompson5910):

no worries, you don't have to rush anything if you go too fast, then you won't learn the best you can

OpenStudy (anonymous):

true!